A teacher wishes to schedule a quiz or exam during part of a two-hour class.The equation 1+e=120 can be used to find the available lecture time "1" given that "e" minutes are planned for the quiz or exam.Determine the ordered pairs that satisfy the equation if the domain is{20,30,45,60}.
Answer by venugopalramana(3286)

A teacher wishes to schedule a quiz or exam during part of a two-hour class.The equation 1+e=120 can be used to find the available lecture time "1" given that "e" minutes are planned for the quiz or exam.Determine the ordered pairs that satisfy the equation if the domain is{20,30,45,60}.
L+E=120
L=120-E
DOMAIN OF E IS (20,30,40,50,60)..FINDING L FROM THE ABOVE EQUATION USING THESE DOMAIN VALUES OF E AND TABULATING,WE GET
E...................20...........30.........40.............50...........60.....
L=120-E.............100..........90.........80.............70...........60....
WRITING AS ORDERED PAIRS ,WE GET
(20,100),(30,90),(40,80),(50,70),(60,60)